e demand for shirts produced by a Canadian manufacturer has been estimated to be P 30 Th - Q/200 a. Compute the point elasticity at P- $10; at P- $15. b. How does the point elasticity vary with the price?
Answer #1

a) P = 30 - Q/200 = 30 - 0.005Q

0.005Q = 30 - P

Q = (30/0.005) - (P / 0.005) = 6000 - 200P

At P = $10

Q = 6000 - 200(10) = 4000

Elasticity = ∆Q/∆P * ( P / Q)                          [Where, ∆Q/∆P = price coefficient in the demand function]

               = -200 * (10 / 4000)

               = -0.5

The absolute value of elasticity is 0.5.

At P = $15

Q = 6000 - 200(15) = 3000

Elasticity = ∆Q/∆P * ( P / Q)                          [Where, ∆Q/∆P = price coefficient in the demand function]

               = -200 * (15 / 3000)

               = -1

The absolute value of elasticity is 1.

b) It is clear from the above calculation that the point elasticity increases as price rises.